log in base 3(2x+1)=2
what does x equal?

Answer by dabanfield(803)

log in base 3(2x+1)=2
what does x equal?
By definition if log base3 (2x+1) = 2 we have:
2x+1 = 3^2
2x+1 = 9
Solve for x.